# Quantitative surgery and total mean curvature

### v1

@@ 5. Surgery bending in arbitrary backgrounds
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\mu})={}$ $u^{-\frac{n+2}{n-2}}\Bigl(\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\frac{4(n-1)}{n-2}(1-\langle\nu_{\varrho},\partial_{t}\rangle^{2})\partial^{2}_{t}u$
$\qquad\qquad+\frac{4(n-1)}{n-2}H(\Sigma_{\varrho},g+\mathrm{d}t^{2})\langle\nu_{\varrho},\partial_{t}\rangle\partial_{t}u\Bigr)$
If $t\geq 2\varrho$, then $\Sigma_{\varrho}$ is given by a cylinder, so $\langle\nu_{\varrho},\partial_{t}\rangle=0$. Furthermore, $u\leq 2$ and $\partial^{2}_{t}u=3\mu\eta^{\prime}(3t-1)+9t\mu\eta^{\prime\prime}(3t-1)\leq 90$ by the choice of $\eta$. Therefore, the above formula gives the following estimate to
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\mu})\geq{}$ $2^{-\frac{n+2}{n-2}}\Bigl(\mathrm{scal}(M,g)+\underbrace{\frac{1}{12}\varrho^{-3}-\frac{360(n-1)}{n-2}}_{\geq 33|\mathrm{scal}(M,g)|\geq 0}\Bigr)\geq|\mathrm{scal}(M,g)|.$
On the other hand, if $t\leq 2\varrho$, then $u=1+t\mu$, $\partial_{t}u=\mu$ and $\partial_{t}^{2}u=0$. As $H(\Sigma_{\varrho},g+\mathrm{d}t^{2})$ is positive, so is the entire final summand, and we get
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\mu})\geq{}$ $2^{-\frac{n+2}{n-2}}\underbrace{\Bigl(\mathrm{scal}(M,g)+\frac{1}{12}\varrho^{-3}\Bigr)}_{\geq 32|\mathrm{scal}(M,g)|\geq 0}\geq|\mathrm{scal}(M,g)|,$
@@
$\mathrm{scal}(\overline{\Sigma}_{\varrho})\geq\frac{1}{3}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}\quad\text{and}\quad\|\nu_{\varrho}^{\intercal}\|\leq\overline{C}\sqrt{\varrho}$
for $\overline{C}=18(n-k)+432$ the constant from Proposition 4.3. Furthermore, we have $\varrho\leq\overline{C}\sqrt{\varrho}$ and therefore, (5.1) from Lemma 5.3 yields
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g)$
$\quad\geq\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})-2C\overline{C}\sqrt{\varrho}\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)^{2}$
$\quad\geq\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})-2C\overline{C}\sqrt{\varrho}-4C\overline{C}\sqrt{\varrho}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|-2C\overline{C}\sqrt{\varrho}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}$
If $|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\leq 1$, then this becomes $\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})\geq\mathrm{scal}(M,g)-8C\overline{C}\sqrt{\varrho}$ since $\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})\geq 0$. On the other hand, if $|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\geq 1$, then $|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\leq|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}$ and we obtain
(5.7) $\begin{split}&\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g)\\ &\qquad\geq\frac{1}{3}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}-2C\overline{C}\sqrt{\varrho}-6C\overline{C}\sqrt{\varrho}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}\\ &\qquad\geq{}-2C\overline{C}\sqrt{\varrho}+\underbrace{\left(\frac{1}{3}-6C\overline{C}\sqrt{\varrho}\right)}_{\geq 0}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}\\ &\qquad\geq{}-2C\overline{C}\sqrt{\varrho},\end{split}$
so $\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})\geq\mathrm{scal}(M,g)-8C\overline{C}\sqrt{\varrho}$. Concerning the mean curvature, Proposition 4.3 yields
$H(\overline{\Sigma}_{\varrho},\delta)\geq\frac{1}{2}|A(\overline{\Sigma}_{\varrho},\delta)|,$
@@
$H(\Sigma_{\varrho},g+\mathrm{d}t^{2})\geq{}$ $H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})-C\overline{C}\sqrt{\varrho}\cdot(2+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|_{\delta+\mathrm{d}t^{2}})$
$\geq{}$ $|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\underbrace{\left(\frac{1}{2}-C\sqrt{\varrho}\right)}_{\geq 0}-2C\overline{C}\sqrt{\varrho}\geq-2C\overline{C}\sqrt{\varrho}$
By employing the conformal change formula from Lemma 5.4, we obtain
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\mu})={}$ $u^{-\frac{n+2}{n-2}}\Bigl(\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})+\frac{4(n-1)}{n-2}H(\Sigma_{\varrho},g+\mathrm{d}t^{2})\langle\nu_{\varrho},\partial_{t}\rangle\mu\Bigr)$
$\geq$ ${}u^{-\frac{n+2}{n-2}}\Bigl(\mathrm{scal}(M,g)-\varrho^{1/4}\underbrace{\left(8C\overline{C}\varrho^{1/4}+C\overline{C}\frac{8(n-1)}{n-2}\varrho^{1/4}\right)}_{\leq\tfrac{1}{2}}\Bigr)$
$\geq$ ${}u^{-\frac{n+2}{n-2}}\left(\mathrm{scal}(M,g)-\frac{1}{2}\varrho^{1/4}\right).$
Since $u=1+t\mu\leq 1+2\varrho\mu$, we have $1\geq u^{-1}\geq u^{-(n+2)/(n-2)}\geq u^{{-5}}\geq 1-5t\mu\geq 1-10\varrho$. If $\mathrm{scal}(M,g)-\frac{1}{2}\varrho^{1/4}\leq 0$, then $\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\mu})\geq\mathrm{scal}(M,g)-\frac{1}{2}\varrho^{1/4}$, so it suffices to consider the case that $\mathrm{scal}(M,g)-\frac{1}{2}\varrho^{1/4}\geq 0$, where we can argue similarly:
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\mu})\geq$ ${}(1-10\varrho)\left(\mathrm{scal}(M,g)-\frac{1}{2}\varrho^{1/4})\right)$
$\geq{}$ $\mathrm{scal}(M,g)-\varrho^{1/4}\underbrace{\Bigl(10|\mathrm{scal}(M,g)|\varrho^{3/4}+\frac{1}{2}\Bigr)}_{\leq 1}\geq\mathrm{scal}(M,g)-\varrho^{1/4}.$
@@
$\|\nu_{\varrho}^{\intercal}\|\leq\overline{C}\sqrt{\varrho},\qquad|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\leq\overline{C}\varrho^{3/8}\quad\text{and}\quad\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho})\geq-\overline{C}\varrho^{3/4}.$
@@
(5.8) $\begin{split}&\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g)\\ &\qquad\geq-\overline{C}\varrho^{3/4}-C\varrho-C\cdot\Bigl(3\overline{C}\varrho^{1/2}+2\overline{C}\varrho^{7/8}+\overline{C}\varrho^{14/8}\Bigr)\\ &\qquad\geq{}\mathrm{scal}(M,g)-8C\overline{C}\sqrt{\varrho}.\end{split}$
Furthermore, we have
$H(\Sigma_{\varrho},g+\mathrm{d}t^{2})\geq{}$ $\underbrace{H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})}_{\geq-n|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|}-C\left(\|\nu_{\varrho}^{\intercal}\|+\varrho^{2}(1+\|\nu_{\varrho}^{\intercal}\|+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)\right)$
$\geq{}$ $-n|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|-2C\sqrt{\varrho}$
$\geq{}$ $-n\varrho^{3/8}-2C\sqrt{\varrho}.$
An application of Lemma 5.4 together with the observation that $\partial_{t}u=\mu$ and $\partial_{t}^{2}u=0$ now gives:
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\mu})\geq{}$ $u^{-\frac{n+2}{n-2}}\Bigl(\mathrm{scal}(M,g)-\varrho^{1/4}\underbrace{\left(8C\overline{C}\varrho^{1/4}+\frac{4(n-1)}{n-2}n\varrho^{1/8}+2C\varrho^{1/4}\right)}_{\leq\frac{1}{2}}\Bigr)$
$\geq{}$ $u^{-\frac{n+2}{n-2}}\left(\mathrm{scal}(M,g)-\frac{1}{2}\varrho^{1/4}\right).$
If $\mathrm{scal}(M,g)-\tfrac{1}{2}\varrho^{1/4}<0$ is negative, then $\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\mu})\geq\mathrm{scal}(M,g)-\tfrac{1}{2}\varrho^{1/4}$, because $u^{-(n+2)/(n-2)}\leq 1$. On the other hand, if $\mathrm{scal}(M,g)-\frac{1}{2}\varrho^{1/4}\geq 0$, we again use $u^{-(n+2)/(n-2)}\geq 1-10\varrho$ to obtain
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\mu})\geq{}$ $(1-10\varrho)(\mathrm{scal}(M,g)-\frac{1}{2}\varrho^{1/4})$
$\geq{}$ $\mathrm{scal}(M,g)-\varrho^{1/4}\underbrace{\left(10|\mathrm{scal}(M,g)|\varrho^{3/4}+\frac{1}{2}\right)}_{\leq 1}$
$\geq{}$ $\mathrm{scal}(M,g)-\varrho^{1/4}$
which finishes the proof of the first half of Theorem 5.1 concerning the scalar curvature of $(\Sigma_{\varrho},g_{\varrho,\mu})$.
It remains to show positivity of the mean curvature with respect to the conformally transformed metric after passing to a smaller radius bound $\varrho_{0}$. To be precise, let
$\varrho_{0}\coloneqq{}$ $\frac{1}{2}\min\left\{\frac{n-2}{104n-16},\,\frac{1}{(3C)^{4}},\,\frac{1}{(n+2C)^{8}},\,\frac{1}{5^{8}},\mu^{8}\right\}$
Let us recall the conformal transformation formula:
$H(\Sigma_{\varrho},g_{\mu})=u^{-\frac{2}{n-2}}\left(H(\Sigma_{\varrho},g+\mathrm{d}t^{2})+\frac{2n}{n-2}\langle\nu_{\varrho},\partial_{t}\rangle u^{-1}\partial_{t}u\right)$
@@
$H(\Sigma_{\varrho},g+\mathrm{d}t^{2})\geq\frac{1}{8}\varrho^{-3/2}$
and, since $\partial_{t}u=\mu\eta(3t-1)+3t\mu\eta^{\prime}(3t-1)\geq-6$, $u^{-1}\leq 1$, $\mu\leq 1$ and $\langle\nu_{\varrho},\partial_{t}\rangle\leq 1$, we thus obtain
$H(\Sigma_{\varrho},g_{\mu})\geq{}$ $u^{-\frac{2}{n-2}}\underbrace{\left(\frac{1}{8}\varrho^{-3/2}-\frac{12n}{n-2}\right)}_{\geq 1}>0,$
which is positive and independent of $\mu$.
For the final case $r\in[\varrho^{2},2\sqrt{\varrho}]$, we first observe, that we either have
$H(g+\mathrm{d}t^{2})\geq-3C\sqrt{\varrho}=-\varrho^{1/4}\underbrace{(3C\varrho^{1/4})}_{\leq 1}\geq-\varrho^{1/4}$
if $r\leq\sqrt{\varrho}$ or
$H(g+\mathrm{d}t^{2})\geq-n\varrho^{3/8}-2C\sqrt{\varrho}=-\varrho^{1/4}\underbrace{(n\varrho^{1/8}-2C\varrho^{1/4})}_{\leq 1}\geq-\varrho^{1/4}$
if $r\geq\sqrt{\varrho}$. Furthermore, we also have
$\langle\nu_{\varrho},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}\geq\langle\overline{\nu}_{\varrho},\partial_{t}\rangle_{\delta+\mathrm{d}t^{2}}-2C\varrho^{2}\geq 1-\sqrt{\varrho}-2C\varrho^{2}\geq\frac{1}{2}.$
We can now plug both of these, together with $\partial_{t}u=\mu$, $1\geq u^{-1}\geq 1-2\varrho\mu$ and $u^{-\frac{2}{n-2}}\geq 1-4\varrho\mu$ into the transformation formula to obtain
$H(\Sigma_{\varrho},g_{\mu})\geq{}$ $u^{-\frac{2}{n-2}}\Bigl(-\varrho^{1/4}+\underbrace{\frac{2n}{n-2}\frac{1}{2}}_{\geq 1}(1-2\varrho\mu)\mu\Bigr)$
$\geq{}$ $(1-4\varrho\mu)\left(\mu-\varrho^{1/4}-\varrho\mu^{2}\right)$
$\geq{}$ $\mu-\varrho^{1/8}(4\mu^{2}\varrho^{7/8}+\varrho^{1/8}).$
Since $\mu^{2}\varrho^{6/8}\leq 1$, we have $(4\mu^{2}\varrho^{7/8}+\varrho^{1/8})\leq 5\varrho^{(1/8)}\leq 1$ and thus
$H(\Sigma_{\varrho},g_{\mu})\geq{}$ $\mu-\varrho^{1/8}$
which is positive by assumption $\varrho_{0}\leq\frac{1}{2}(\mu)^{8}$ and thus finishes the proof of Theorem 5.1. ∎
Let us now delve into the proof of Lemma 5.3.

### v2

@@ 6. Surgery bending in arbitrary backgrounds
$\begin{split}\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\kappa})={}&u^{-\frac{n+2}{n-2}}\Bigl(\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\frac{4(n-1)}{n-2}(1-\langle\nu_{\varrho},\partial_{t}\rangle^{2})\partial^{2}_{t}u\\ &\qquad\qquad+\frac{4(n-1)}{n-2}H(\Sigma_{\varrho},g+\mathrm{d}t^{2})\langle\nu_{\varrho},\partial_{t}\rangle\partial_{t}u\Bigr)\end{split}$
If $t\geq 2\varrho$, then $\Sigma_{\varrho}$ is given by the straight cylinder of radius $\varrho^{4}$, so $\langle\nu_{\varrho},\partial_{t}\rangle=0$. Furthermore, $u\leq 2$ and
$\partial^{2}_{t}u=\alpha^{-1}\tilde{\kappa}\eta^{\prime}(\alpha^{-1}t-1)+\alpha^{-2}t\tilde{\kappa}\eta^{\prime\prime}(\alpha^{-1}t-1)\lesssim\alpha^{-2}\varrho_{1}\lesssim\varrho_{1}^{-1},$
by the choice of $\eta$ and $\kappa$. Therefore, we may assume $\partial^{2}_{t}u\leq C_{1}\varrho_{1}$ and the above formula gives the following estimate
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\kappa})>{}$ $2^{-\frac{n+2}{n-2}}\Bigl(\mathrm{scal}(M,g_{M})+\varrho^{-3}(c_{1}-\frac{4(n-1)}{n-2}C_{0}\varrho_{1}^{2})\Bigr).$
We choose $\varrho_{1}$ such that $c_{1}-\tfrac{4(n-1)}{n-2}C_{0}\varrho_{1}^{2}\leq c_{1}/2$ and $c_{1}\varrho_{1}^{-3/2}\geq 66\max|\mathrm{scal}(M,g_{M})|$, so that we get $\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\kappa})>|\mathrm{scal}(M,g_{M})|$ as desired.
On the other hand, if $t\leq 2\varrho\leq\alpha$, then $u=1+t\tilde{\kappa}$, $\partial_{t}u=\tilde{\kappa}$ and $\partial_{t}^{2}u=0$. As $H(\Sigma_{\varrho},g+\mathrm{d}t^{2})$ is positive, so is the entire final summand, and we get
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\kappa})>{}$ $2^{-\frac{n+2}{n-2}}\underbrace{\Bigl(\mathrm{scal}(M,g_{M})+c_{1}\varrho^{-3}\Bigr)}_{\geq 32\max|\mathrm{scal}(M,g_{M})|}\geq|\mathrm{scal}(M,g_{M})|,$
@@
$\mathrm{scal}(\overline{\Sigma}_{\varrho})>\frac{1}{3}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}\quad\text{and}\quad\|\nu_{\varrho}^{\intercal}\|<\overline{C}\sqrt{\varrho}$
for $\overline{C}=18k+432$ the constant from Proposition 5.3. We set $C_{2}\coloneqq 4C_{0}\cdot\overline{C}$, and observe that (6.6) yields
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g_{M})$
$\qquad>\frac{1}{3}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}$
$\qquad\qquad-\frac{C_{2}}{4}\left(\varrho\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)^{2}-\varrho^{1/2}\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)\right)$
$\qquad\geq-\frac{C_{2}}{4}\varrho^{1/2}-\frac{C_{2}}{2}\varrho^{1/2}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|+\left(\frac{1}{3}-\frac{C_{2}}{4}\varrho\right)|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}$
If $|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\leq 1$, then this becomes $\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g_{M})>-C_{2}\varrho^{1/2}$. On the other hand, if $|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\geq 1$, then $|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\leq|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}$ and we obtain
$\begin{split}&\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g_{M})>-\frac{C_{2}}{4}\varrho^{1/2}+\left(\frac{1}{3}-\frac{C_{2}}{2}\varrho^{1/2}-\frac{C_{2}}{4}\varrho\right)|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|^{2}\end{split}$
and thus, for $\varrho_{1}\leq(\frac{1}{3C_{2}})^{2}$, we have
(6.10) $\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})>\mathrm{scal}(M,g_{M})-C_{2}\varrho^{1/2}$
Concerning the mean curvature, Proposition 5.3 again yields
$H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})>\frac{1}{2}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|,$
@@
$\begin{split}H(\Sigma_{\varrho},g+\mathrm{d}t^{2})>{}&\frac{1}{2}|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|-C_{1}\cdot\overline{C}\left(\varrho^{1/2}+\varrho\cdot(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)\right)\\ \geq{}&-C_{3}\varrho^{1/2}+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\left(\frac{1}{2}-C_{3}\varrho\right)\end{split}$
where $C_{3}\coloneqq 2C_{1}\cdot\overline{C}$. If $\varrho_{1}\leq\tfrac{1}{4C_{3}}$, we get
(6.11) $H(\Sigma_{\varrho},g+\mathrm{d}t^{2})>-C_{3}\varrho^{1/2}$
By employing the conformal change formula from Lemma 6.5, we can use $u=1+t\tilde{\kappa}$ to obtain
[…]
$>{}$ $u^{-\frac{n+2}{n-2}}\Bigl(\mathrm{scal}(M,g_{M})-C_{3}\varrho^{1/2}\Bigr)$
$\geq{}$ $u^{-\frac{n+2}{n-2}}\left(\mathrm{scal}(M,g_{M})-\frac{1}{2}\varrho^{1/4}\right),$
if $\varrho_{1}^{1/4}C_{3}\leq\tfrac{1}{2}$. Since $u=1+t\tilde{\kappa}\leq 1+\kappa\leq 1+\varrho_{1}$, we have $1\geq u^{-1}\geq u^{-(n+2)/(n-2)}\geq u^{{-5}}\geq 1-5t\tilde{\kappa}\geq 1-5\varrho_{1}$. [^2] This follows from the inequality $(1+x)^{-n}\geq 1-nx$ for $n\in\mathbb{N}$. If $\mathrm{scal}(M,g_{M})-\frac{1}{2}\varrho^{1/4}\leq 0$, then
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\kappa})>\mathrm{scal}(M,g_{M})-\frac{1}{2}\varrho^{1/4}\geq\mathrm{scal}(M,g_{M})-\frac{1}{2}\varrho_{1}^{1/4},$
so it suffices to consider the case that $\mathrm{scal}(M,g_{M})-\frac{1}{2}\varrho^{1/4}\geq 0$, where we can argue similarly:
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\kappa})>{}$ $(1-5\varrho_{1})\left(\mathrm{scal}(M,g_{M})-\frac{1}{2}\varrho^{1/4}\right)$
$>{}$ $\mathrm{scal}(M,g_{M})-\varrho_{1}^{1/4}\Bigl(5\cdot\mathrm{scal}(M,g_{M})\varrho_{1}^{3/4}+\frac{1}{2}\Bigr),$
and, if $\varrho_{1}$ is chosen such that $5\max\mathrm{scal}(M,g_{M})\varrho_{1}^{3/4}<\frac{1}{2}$, we obtain
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\kappa})>\mathrm{scal}(M,g_{M})-\varrho_{1}^{1/4}.$
@@
$\|\nu_{\varrho}^{\intercal}\|\lesssim\varrho^{1/2},\qquad|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\lesssim\varrho^{3/8}\quad\text{and}\quad\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho})\gtrsim-\varrho^{3/4}.$
@@
(6.12) $\begin{split}&\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g_{M})\\ &\qquad\gtrsim-\varrho^{3/4}-\varrho\cdot(1+\varrho^{3/8})^{2}-\varrho^{1/2}\cdot(1+\varrho^{3/8})\gtrsim-\varrho^{1/2}\end{split}$
Furthermore, using $H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})\geq-n|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|\gtrsim\varrho^{3/8}$ we obtain
(6.13) $\begin{split}H(\Sigma_{\varrho},g+\mathrm{d}t^{2})\gtrsim{}&H(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})-\|\nu_{\varrho}^{\intercal}\|-\varrho(1+|A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})|)\\ \gtrsim{}&-\varrho^{3/8}-\varrho^{1/2}-\varrho-\varrho^{11/8}\gtrsim-\varrho^{3/8}\end{split}$
Thus there exists a constant $C_{4}>0$ such that
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})>\mathrm{scal}(M,g_{M})-C_{4}\varrho^{1/2}\quad\text{and}\quad H(\Sigma_{\varrho},g+\mathrm{d}t^{2})>-C_{4}\varrho^{3/8}.$
An application of Lemma 6.5 together with the observation that $\partial_{t}u=\tilde{\kappa}\leq 1$, $\partial_{t}^{2}u=0$ and $\tfrac{4(n-1)}{(n-2)}\leq 8$ now gives:
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\kappa})>{}$ $u^{-\frac{n+2}{n-2}}\left(\mathrm{scal}(M,g_{M})-\varrho^{1/4}C_{4}\cdot(\varrho^{1/4}+8\varrho^{1/8})\right)$
$\geq{}$ $u^{-\frac{n+2}{n-2}}\left(\mathrm{scal}(M,g_{M})-\frac{1}{2}\varrho^{1/4}\right),$
provided $\varrho_{1}$ satisfies $C_{4}\cdot(\varrho^{1/4}+8\varrho^{1/8})<\tfrac{1}{2}$. Using $\varrho^{3/8}\leq\varrho^{1/4}$, the same argument as in the second case, below (6.11) shows that
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho,\kappa})>\mathrm{scal}(M,g_{M})-\varrho_{1}^{1/4}.$
It remains to show positivity of the mean curvature with respect to the conformally transformed metric after passing to a smaller radius bound $\varrho_{0}$ which may depend on $\kappa$. Recall the conformal transformation formula:
$H(\Sigma_{\varrho},g_{\kappa})=u^{-\frac{2}{n-2}}\left(H(\Sigma_{\varrho},g+\mathrm{d}t^{2})+\frac{2n}{n-2}\langle\nu_{\varrho},\partial_{t}\rangle u^{-1}\partial_{t}u\right)$
@@
$H(\Sigma_{\varrho},g+\mathrm{d}t^{2})>c_{1}\varrho^{-3/2},$
see (6.9) Since $\partial_{t}u=\tilde{\kappa}\eta(\alpha^{-1}t-1)+\alpha^{-1}t\tilde{\kappa}\eta^{\prime}(\alpha^{-1}t-1)\geq-3\alpha^{-1}$, $u^{-1}\leq 1$, $\tilde{\kappa}\leq 1$ and $\langle\nu_{\varrho},\partial_{t}\rangle\leq 1$, we thus obtain for $C_{5}\coloneqq\tfrac{6n\alpha^{-1}}{(n-2)}$:
(6.14) $H(\Sigma_{\varrho},g_{\kappa})>u^{-\frac{2}{n-2}}\left(c_{1}\varrho^{-3/2}-C_{5}\right).$
If $\varrho_{0}$ is chosen such that $c_{1}\varrho_{0}^{-3/2}-C_{5}\geq 2^{2/(n-2)}\geq u^{2/(n-2)}\kappa$, then
$H(\Sigma_{\varrho},g_{\kappa})>\kappa>(1-\alpha)\kappa-\varrho^{1/4}$
as desired.
The cases $r\in[\varrho^{2},\sqrt{\varrho}]$ and $r\in[\sqrt{\varrho},2\sqrt{\varrho}]$ can be treated simultaneously. We first observe, that (6.11) and (6.13) both give
(6.15) $H(\Sigma_{\varrho},g+\mathrm{d}t^{2})>-\varrho^{1/4},$
provided $\varrho_{0}$ is small enough. Furthermore, combining $\langle\nu_{\varrho},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}-\langle\overline{\nu}_{\varrho},\partial_{t}\rangle_{\delta+\mathrm{d}t^{2}}\gtrsim-\varrho$ from Lemma 6.4 and $\langle\overline{\nu}_{\varrho},\partial_{t}\rangle_{\delta+\mathrm{d}t^{2}}-1\gtrsim-\varrho^{1/2}$ from Proposition 5.3, we find a constant $C_{6}>0$ such that we have
(6.16) $\langle\nu_{\varrho},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}\geq 1-C_{6}\varrho^{1/2}\geq 1-\frac{1}{2}\varrho^{1/4}$
if $\varrho_{0}\leq C_{6}^{-4}$. From (6.15) and (6.16), together with $\partial_{t}u=\tilde{\kappa}=\tfrac{n-2}{2n}\kappa$ and $u^{-1}\geq 1-\varrho_{1}$, we obtain
$H(\Sigma_{\varrho},g+\mathrm{d}t^{2})+\frac{2n}{n-2}\langle\nu_{\varrho},\partial_{t}\rangle u^{-1}\partial_{t}u$ $>-\frac{1}{2}\varrho^{1/4}+(1-\varrho^{1/4})(1-\varrho_{1})\kappa$
$=(1-\varrho_{1})\kappa-\frac{1}{2}\varrho^{1/4}-\varrho^{1/4}(1-\varrho_{1})\kappa$
$\geq(1-\varrho_{1})\kappa-\varrho^{1/4}$
Note, that $u^{-\frac{2}{n-2}}\geq 1-2\varrho_{1}$, and we choose $\varrho_{0}\leq\tfrac{1}{2}(1-\varrho_{1})^{4}\kappa^{4}$ small enough, so that $(1-\varrho_{1})\kappa-\varrho^{1/4}>0$ as well. The transformation formula (6.5) then gives
(6.17) $\begin{split}H(\Sigma_{\varrho},g_{\kappa})>{}&(1-2\varrho_{1})\Bigl((1-\varrho_{1})\kappa-\varrho^{1/4}\Bigr)\\ ={}&(1-3\varrho_{1})\kappa-\varrho^{1/4}\geq(1-\alpha)\kappa-\varrho^{1/4},\end{split}$
which finishes the proof. ∎
Let us now dive into the proof of Lemma 6.4.